# Anticipatory Music Transformer

### v1

@@ Abstract
We introduce anticipation: a method for constructing a controllable generative model of a temporal point process (the event process) conditioned asynchronously on realizations of a second, correlated process (the control process). We achieve this by interleaving sequences of events and controls, such that controls appear following stopping times in the event sequence. This work is motivated by problems arising in the control of symbolic music generation. We focus on infilling control tasks, whereby the controls are a subset of the events themselves, and conditional generation completes a sequence of events given the fixed control events. We train anticipatory infilling models using the large and diverse Lakh MIDI music dataset. These models match the performance of autoregressive models for prompted music generation, with the additional capability to perform infilling control tasks, including accompaniment. Human evaluators report that an anticipatory model produces accompaniments with similar musicality to even music composed by humans over a $20$-second clip.
@@ 1 Introduction
Imagine you are given a melody, and asked to compose a harmonizing accompaniment. This melody is an example of a temporal point process: a sequence of (musical) events that arrive stochastically at points in time. An accompaniment to this melody is also a temporal point process. The events in the accompaniment are tightly correlated—but often asynchronous—with the events in the melody. Generating an accompaniment to a given melody is an example of a control task: we seek the ability to generate an accompaniment (the events) conditioned on a given melody (the controls). Models that generate symbolic music (i.e., compose) subject to user-specified controls are of broad interest as tools for music co-creation (Louie et al., 2020).
@@
When the events and controls are asynchronous, simple approaches to interleaving these sequences make sampling from the ensuing joint model intractable. This includes the natural sort order, whereby we interleave a control on time $t$ as if it were at time $t-\delta$. For sampling to be tractable, we will see in Section 3 that the index in the interleaved sequence that immediately precedes a control must be a stopping time (Billingsley, 1995). We propose a method for interleaving asynchronous sequences of events and controls such that a control on time $t$ appears in the interleaved sequence at a stopping time close to events near time $t-\delta$. We call this method anticipation. The interval $\delta>0$ is a hyperparameter chosen to be long enough to give the model time to account for (i.e. anticipate) upcoming controls, but short enough to maintain proximity of events and controls (if $\delta=\infty$, we recover Seq2Seq modeling). The interleaved structure of anticipation is visualized in Figure 1.
Figure 1: We construct conditional generative models for sequences of events $\mathbf{e}_{1:N}$, conditioned on controls $\mathbf{u}_{1:K}$. We serialize these paired sequences to define an autoregressive factorization of the joint distribution over events and controls. Anticipation interleaves event and control sequences so that a control $\mathbf{u}_{k}$ on time $s_{k}$ appears in the recent history when predicting events near time $s_{k}$. An anticipated control $\mathbf{u}_{k}$ on time $s_{k}$ appears as if it were at approximately time $s_{k}^{\prime}=s_{k}-\delta$. For example, when predicting $\mathbf{e}_{j+7}$ above, the recent context of the anticipation sequence contains contains past events and controls, as well as the future control $\mathbf{u}_{k+5}$; we say that a model predicting $\mathbf{e}_{j+7}$ given this context anticipates the control $\mathbf{u}_{k+5}$, approximately $\delta$ seconds in advance. Crucially, to be able to condition on controls, the index that immediately preceeds each control in the serialized sequence must be a stopping time, a property that naively interleaving events and controls using the sort order of times $s_{k}^{\prime}$ does not satisfy.
@@
We define an arrival-time encoding of temporal point processes that is amenable to expressive autoregressive sequence modeling and facilitates anticipation (Section 2). We describe the interleaved structure of an anticipatory autoregressive model, and how to train and sample from this model (Section 3). We apply anticipation to construct anticipatory infilling models for music, trained on the Lakh MIDI music dataset Raffel (2016). These models unlock new control capabilities for music generation without sacrificing the performance of unconditional generation (Section 4). We release all code for reproducing these models, along with pre-trained model weights. [^1] For assets and supplemental material, see: https://johnthickstun.com/anticipation/
@@ 2 Music as a Temporal Point Process
A marked temporal point process consists of sparse events situated at points in continuous time (Daley and Vere-Jones, 2007).
@@
A marked temporal point process consists of time-ordered events $\mathbf{e}_{i}=(t_{i},\mathbf{m}_{i})$, where $t_{i}\in\mathbb{R}_{+}$ ($t_{i}\leq t_{j}$ if $i<j$) is a point in time and $\mathbf{m}_{i}\in\mathcal{V}$ is a mark from a finite vocabulary $\mathcal{V}$.
Given controls $\mathbf{u}_{1:K}$ provided by a user, we say that we can control generation of the event process $\mathbf{e}_{1:N}$ with respect to $\mathbf{u}_{1:K}$ if we can sample from $p(\mathbf{e}_{1:N}|\mathbf{u}_{1:K})$. We focus on infilling control, whereby the controls $\mathbf{u}_{1:K}$ share a vocabulary with the events $\mathbf{e}_{1:N}$. Given a user-specified set of $K$ events $\mathbf{u}_{1:K}$, we would like to generate a complete realization of the process $\mathbf{e}_{1:N}$ such that $\mathbf{u}_{1:K}\subseteq\mathbf{e}_{1:N}$. This generalizes the span-infilling task—which ask us to generate a missing contiguous span of events—previously studied in the music literature (Ippolito et al., 2018; Chang et al., 2021; Tan et al., 2021, 2022).
Music can be described as a marked temporal point process, where the vocabulary of marks consists of musical notes and other musical events. In this work, we restrict the vocabulary of musical events to notes consisting of a pitch $\mathbf{p}$, instrument class $\mathbf{k}$, and duration $\mathbf{d}$. Following the MIDI standard (International MIDI Association, 1988), we quantize pitch $\mathbf{p}\in\{0,\dots,127\}$ according to the $12$-tone Western scale ($p=60$ denotes “middle C”, or $261$Hz); we represent instrument classes $\mathbf{k}\in\{0,\dots,128\}$ using standard MIDI instrument codes, including drums (code $128$); we represent duration in units of seconds, quantized to $10$ms intervals with a maximum duration of $10$ seconds. We represent combined pitch and instrument information using a single value $\mathbf{n}=128\mathbf{k}+\mathbf{p}$. A mark is thus a note $\mathbf{m}_{i}=(\mathbf{d}_{i},\mathbf{n}_{i})\in\mathcal{V}$ from a vocabulary of $17,512$ marks, and music consists of these notes situated at points in time.
@@
Stochastic arrival times $t_{i}$ are the defining characteristic of a temporal point process. In Section 2.1 we propose an arrival-time tokenization that represents a realization of a marked point process as a sequence of (discretized) arrival times and marks. In Section 3 we exploit an invariance of this sequence representation (context-free subsequences) to create anticipation. We also consider an interarrival-time tokenization that is comparable to encodings used in recent work on music generation (Huang et al., 2018; Donahue et al., 2019; Payne, 2019; Oore et al., 2020). Both of these encodings reduce point process modeling to sequence density estimation. This allows us to directly apply the full modern machinery of causal autoregressive transformers and large language models to modeling point process data. However, only the proposed arrival-time encoding facilitates the construction of anticipation.
Related Work on Modeling Point Processes.
In the temporal point processes literature, substantial work focuses on modeling a conditional intensity function (Du et al., 2016; Omi et al., 2019; Zuo et al., 2020), e.g., the Hawkes process (Mei and Eisner, 2017). In contrast, we directly model the point process itself (density estimation). In this regard, our work is most comparable to Shchur et al. (2020); but whereas that work models continuous densities, we model discrete distributions over quantized time values. Older work on music rasterizes time with a uniformly sampled grid (Boulanger-Lewandowski et al., 2012) (i.e., a piano-roll). In the piano-roll setting, there are strong solutions to accompaniment and infilling tasks that avoid the complexity of asynchronicity that we address in this paper (Hadjeres et al., 2017; Huang et al., 2017; Pati et al., 2019; Chen et al., 2020). However, for diverse and rhythmically intricate music data (including Lakh MIDI) piano-roll rasterization comes at a cost: if the rasterization is coarse then rhythmic detail is lost, and if the rasterization is fine then the dimensionality of these piano-rolls becomes unwieldy.
@@ 2.1 Encoding Music as Sequences
The more common practice in symbolic music modeling encodes music as a sequence of onset and offset events, separated by interarrival times $\Delta\in\mathbb{R}_{+}$. For each event $\mathbf{e}=(\mathbf{t},\mathbf{m})$, we define onset $\textbf{on}=(\mathbf{t},\mathbf{n})$ and offset $\textbf{off}=(\mathbf{t}+\mathbf{d},\mathbf{n})$.
@@
In contrast to Definition 2.2, the interarrival-time tokenization described by Definition 2.3—and other common music encodings including including REMI (Huang and Yang, 2020) and OctupleMidi (Zeng et al., 2021)—are context sensitive: the timings of tokens are determined contextually by their positions in the sequence. Interarrival-time tokenization may appear less compact than arrival-time tokenization: each event is encoded as an onset and an offset—each with an associated interarrival time—totaling $4$ tokens, versus a triplet of arrival time, duration, and note totaling $3$ tokens. However, omitting zero-duration interrarival times results in sequences of comparable length under either encoding for music datasets including Lakh MIDI.
@@ 3 Anticipation
$p(\mathbf{a}_{1:N})=\prod_{i=1}^{N+K}p(\mathbf{a}_{i}|\mathbf{a}_{1:i-1}).$ (5)
@@
Therefore $\mathbf{a}_{1:4}=(\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{u}_{1},\mathbf{e}_{3})$. Contrast this order with the more natural sort order that interleaves $\mathbf{u}_{1}$ as if it were an event at time $s_{1}-\delta=2$: $(\mathbf{e}_{1},\mathbf{u}_{1},\mathbf{e}_{2},\mathbf{e}_{3})$. Under the sort order, $\mathbf{u}_{1}$ appears before the first event $\mathbf{e}_{j}$ exceeding time $s_{1}-\delta$. This position can only be determined retroactively (after observing the position of $\mathbf{e}_{2}$) which precludes autoregressive inference. Formally, the problem here is that $\tau_{\mathbf{u}_{1}}-1$ is not a stopping time.
@@ 3.2 Sparse Sequences
In this case, the anticipatory interleaving of $\mathbf{u}_{1}$ with $\mathbf{e}^{\prime}_{1:5}$ is $\mathbf{a}^{\prime}=(\mathbf{e}^{\prime}_{1},\mathbf{e}^{\prime}_{2},\mathbf{e}^{\prime}_{3},\mathbf{u}_{1},\mathbf{e}^{\prime}_{4},\mathbf{e}^{\prime}_{5})$.
And in this case, we anticipate the control $\mathbf{u}_{1}$ on time $s_{1}=4.5$ between the event $\mathbf{e}^{\prime}_{3}$ at time $3$ and the event $\mathbf{e}^{\prime}_{4}$ at time $4$. All the anticipatory models in this paper are trained with $\Delta^{*}=1$ second.
@@ 3.3 Training Anticipatory Models
We train anticipatory autoregressive models using standard maximum likelihood estimation of the sequences $\mathbf{a}_{1:N+K}$ (Definition 3.1). We tokenize these sequences according to the encoding described in Definition 2.2: $\mathbf{x}_{1:3(N+K)}\equiv\texttt{tokenize}(\mathbf{a}_{1:N+K})$. We follow a standard sequence packing procedure (see, e.g, Appendix B of Brown et al. (2020)) to construct training examples of fixed length $M$ (the model context) from variable-length sequences $\mathbf{x}_{1:3(N+K)}$, using a special event SEP as the sequence separator. We prepend each training example with a single global control code $\mathbf{z}\in\{0,1\}$ that indicates whether the example contains local controls $\mathbf{u}_{k}$; setting $\mathbf{z}=0$ (no controls) facilitates comparisons between anticipatory and autoregressive models. If the training example spans multiple sequences, then $\mathbf{z}$ describes the sequence preceding the first SEP event in the example. We randomly shuffle and mini-batch training examples for stochastic gradient training.
@@
Dividing up a dataset into training examples of length $M$ introduces boundary effects. For general autoregressive models, this procedure results in an “early token curse,” whereby predictions early in a training example must be made with limited context (Press et al., 2021). For anticipatory autoregressive models, more subtle boundary effects arise. For an anticipation interval $\delta$, controls on first $\delta$ seconds of the training example do not appear in the context: these controls appear at the end of the previous training example. If $\delta$ is large relative to $M$ then the boundary effects become severe: many event predictions will be made without the relevant contextual controls, and vice-versa. This tempers the value of making predictions for all $M-1$ sequence indices, imposing a drag on training efficiency for large values $\delta$ relative to $M$. Like the early token curse, this effect is mitigated with larger model contexts $M$ relative to $\delta$; in practice, the maximum practical anticipation interval $\delta$ is thus coupled with the context size of the model that we plan to train.
@@
$\mathbf{a}_{0}\leftarrow\texttt{SEP}$   // A special sequence separator event
$i\leftarrow 1$   // Index $i$ tracks position in the generated sequence
$k\leftarrow 0$   // Index $k$ tracks position in the control sequence
@@
       $t\leftarrow\texttt{time}(\mathbf{a}_{i-1})$   // Get the time $t$ of the previous event $\mathbf{a}_{i-1}$
       while $\emph{{time}}(\mathbf{u}_{k})\leq t+\delta$   // While there are controls before time $t+\delta$
@@
             $\mathbf{a}_{i}\leftarrow\mathbf{u}_{k}$   // Anticipate control $\mathbf{u}_{k}$ at index $i$
             $i\leftarrow i+1$   // Advance to index $i+1$
             $k\leftarrow k+1$   // Consume control $\mathbf{u}_{k}$
@@
      Sample $\mathbf{a}_{i}\sim p(\cdot|\mathbf{z},\mathbf{a}_{i-M-\text{Length}(\mathbf{z}):i-1})$   // Sample an event from the model
       $i\leftarrow i+1$   // Advance to index $i+1$

### v2

@@ Abstract
We introduce anticipation: a method for constructing a controllable generative model of a temporal point process (the event process) conditioned asynchronously on realizations of a second, correlated process (the control process). We achieve this by interleaving sequences of events and controls, such that controls appear following stopping times in the event sequence. This work is motivated by problems arising in the control of symbolic music generation. We focus on infilling control tasks, whereby the controls are a subset of the events themselves, and conditional generation completes a sequence of events given the fixed control events. We train anticipatory infilling models using the large and diverse Lakh MIDI music dataset. These models match the performance of autoregressive models for prompted generation, with the additional capability to perform infilling control tasks, including accompaniment. Human evaluators report that an anticipatory model produces accompaniments with similar musicality to even music composed by humans over a $20$-second clip.
@@ 1 Introduction
Imagine you are given a melody, and asked to compose a harmonizing accompaniment. This melody can be modeled by a temporal point process: a probability distribution over (musical) events that arrive stochastically at points in time. An accompaniment to this melody is also a realization of a temporal point process. The events in the accompaniment are tightly correlated—but often asynchronous—with the events in the melody. Generating an accompaniment to a given melody is an example of a control task: we seek the ability to generate an accompaniment (the events) conditioned on a given melody (the controls). Models that generate symbolic music (i.e., compose) subject to user-specified controls are of broad interest as tools for music co-creation (Louie et al., 2020).
@@
When events and controls are asynchronous, simple approaches to interleaving these sequences make sampling from the ensuing joint model intractable. This includes the natural sort order, whereby we interleave a control on time $t$ as if it were at time $t-\delta$. For tractable sampling, we will see in Section 3 that the index in the interleaved sequence that immediately precedes a control must be a stopping time (Billingsley, 1995). We propose a method for interleaving asynchronous events and controls such that a control on time $t$ appears in the interleaved sequence at a stopping time close to events near time $t-\delta$. We call this method anticipation. The interval $\delta>0$ is a hyperparameter chosen to be long enough to give the model time to account for (i.e. anticipate) upcoming controls, but short enough to maintain proximity of events and controls (if $\delta=\infty$, we recover Seq2Seq modeling). The interleaved structure of anticipation is visualized in Figure 1.
Figure 1: We construct generative models for sequences of events $\mathbf{e}_{1:N}$, conditioned on controls $\mathbf{u}_{1:K}$. We serialize these paired sequences to define an autoregressive factorization of the joint distribution over events and controls. Anticipation interleaves event and control sequences so that a control $\mathbf{u}_{k}$ on time $s_{k}$ appears in the recent history when predicting events near time $s_{k}$. An anticipated control $\mathbf{u}_{k}$ on time $s_{k}$ appears as if it were at approximately time $s_{k}^{\prime}=s_{k}-\delta$. For example, when predicting $\mathbf{e}_{j+7}$ above, the recent context of the anticipation sequence contains contains past events and controls, as well as the future control $\mathbf{u}_{k+5}$; we say that a model predicting $\mathbf{e}_{j+7}$ given this context anticipates the control $\mathbf{u}_{k+5}$, approximately $\delta$ seconds in advance. Crucially, to be able to condition on controls, the index that immediately preceeds each control in the serialized sequence must be a stopping time, a property that naively interleaving events and controls using the sort order of times $s_{k}^{\prime}$ does not satisfy.
@@
We define an arrival-time encoding of events and controls that is amenable to expressive autoregressive sequence modeling and facilitates anticipation (Section 2). We describe the interleaved structure of an anticipatory autoregressive model, and how to train and sample from this model (Section 3). We apply anticipation to construct anticipatory infilling models for music, trained on the Lakh MIDI music dataset (Raffel, 2016). These models unlock new control capabilities for music generation without sacrificing the performance of unconditional generation (Section 4). We release all code for reproducing these models, along with pre-trained model weights. [^1] For assets and supplemental material, see: https://johnthickstun.com/anticipation/
@@ 2 Music as a Temporal Point Process
A marked temporal point process is a probability distribution over sparse events situated at points in continuous time (Daley & Vere-Jones, 2007).
@@
A marked temporal point process is a probability distribution over events $\mathbf{e}_{i}=(t_{i},\mathbf{m}_{i})$, where $t_{i}\in\mathbb{R}_{+}$ ($t_{i}\leq t_{j}$ if $i<j$) is a point in time and $\mathbf{m}_{i}\in\mathcal{V}$ is a mark from a finite vocabulary $\mathcal{V}$.
Given controls $\mathbf{u}_{1:K}$ provided by a user, we say that we can control generation of the events $\mathbf{e}_{1:N}$ with respect to $\mathbf{u}_{1:K}$ if we can sample from $p(\mathbf{e}_{1:N}|\mathbf{u}_{1:K})$. We focus on infilling control, whereby the controls $\mathbf{u}_{1:K}$ share a vocabulary with the events $\mathbf{e}_{1:N}$. Given a user-specified set of $K$ events $\mathbf{u}_{1:K}$, we would like to generate a complete realization of the process $\mathbf{e}_{1:N}$ such that $\mathbf{u}_{1:K}\subseteq\mathbf{e}_{1:N}$. This generalizes the span-infilling task—which asks us to generate a missing contiguous span of events—previously studied in the music literature (Ippolito et al., 2018; Chang et al., 2021; Tan et al., 2022a; b).
Music can be described by a marked temporal point process, where the vocabulary of marks consists of musical notes and other musical events. In this work, we restrict the vocabulary of musical events to notes consisting of a pitch $\mathbf{p}$, instrument class $\mathbf{k}$, and duration $\mathbf{d}$. Following the MIDI standard (International MIDI Association, 1988), we quantize pitch $\mathbf{p}\in\{0,\dots,127\}$ according to the $12$-tone Western scale ($p=60$ denotes “middle C”, or $261$Hz); we represent instrument classes $\mathbf{k}\in\{0,\dots,128\}$ using standard MIDI instrument codes, including drums (code $128$); we represent duration in units of seconds, quantized to $10$ms intervals with a maximum duration of $10$ seconds. We represent combined pitch and instrument information using a single value $\mathbf{n}=128\mathbf{k}+\mathbf{p}$. A mark is thus a note $\mathbf{m}_{i}=(\mathbf{d}_{i},\mathbf{n}_{i})\in\mathcal{V}$ from a vocabulary of $17,512$ marks, and music consists of these notes situated at points in time.
Older work on music generation typically rasterizes time, encoding music as a uniformly sampled matrix or tensor, i.e., a piano-roll (Boulanger-Lewandowski et al., 2012; Dong et al., 2018). In the piano-roll setting, there are strong solutions to accompaniment and infilling tasks that avoid the complexities of asynchronicity that we address in this paper (Hadjeres et al., 2017; Huang et al., 2017; Pati et al., 2019; Chen et al., 2020). However, for diverse and rhythmically intricate music data (including the Lakh MIDI dataset) piano-roll rasterization comes at a cost: if the rasterization is coarse then rhythmic detail is lost, and if the rasterization is fine then the piano-rolls are high-dimensional, presenting computational challenges. We instead chose to model music as a temporal point process, an approach inspired by Huang et al. (2018); Oore et al. (2020).
Modeling Temporal Point Processes.
Much of the machine learning literature on temporal point processes focuses on modeling conditional intensity functions (Du et al., 2016; Mei & Eisner, 2017; Omi et al., 2019; Zuo et al., 2020). In contrast, we will model the probability distribution over the next event in a time-ordered sequence. In this regard, our work is most comparable to Shchur et al. (2020); but whereas that work models continuous densities, we model discrete distributions over quantized time values. This allows us to directly apply the full modern machinery of causal autoregressive transformers and large language models to modeling point processes.
@@
Stochastic arrival times $t_{i}$ are the defining characteristic of a temporal point process. In Section 2.1 we propose an arrival-time tokenization that represents a realization of a marked point process as a sequence of (discretized) arrival times and marks. In Section 3 we exploit an invariance of this sequence representation (context-free subsequences) to create anticipation. We also consider an interarrival-time tokenization that is comparable to encodings used in recent work on music generation (Huang et al., 2018; Donahue et al., 2019; Payne, 2019; Oore et al., 2020). Both of these encodings reduce point process modeling to sequence density estimation. However, only the proposed arrival-time encoding facilitates the construction of anticipation.
@@ 2.1 Encoding Music as Sequences
The more common practice in symbolic music modeling encodes music as a sequence of onset and offset events, separated by interarrival times $\Delta\in\mathbb{R}_{+}$. For each event $\mathbf{e}=(\mathbf{t},\mathbf{m})$, we define onset $\textbf{on}=(\mathbf{t},\mathbf{n})$ and offset $\textbf{off}=(\mathbf{t}+\mathbf{d},\mathbf{n})$. We bound the interarrival time between events by 10 seconds, corresponding to a vocabulary of 1000 possible (10ms quantized) interarrival times.
@@
In contrast to Definition 2.2, the interarrival-time tokenization described by Definition 2.3—and other common music encodings including including REMI (Huang & Yang, 2020) and OctupleMidi (Zeng et al., 2021)—are context sensitive: the timings of tokens are determined contextually by their positions in the sequence. Interarrival-time tokenization may appear less compact than arrival-time tokenization: each event is encoded as an onset and an offset—each with an associated interarrival time—totaling $4$ tokens, versus a triplet of arrival time, duration, and note totaling $3$ tokens. However, omitting zero-duration interrarival times results in sequences of comparable length under either encoding of the Lakh MIDI dataset.
@@ 3 Anticipation
$p(\mathbf{a}_{1:N+K})=\prod_{i=1}^{N+K}p(\mathbf{a}_{i}|\mathbf{a}_{1:i-1}).$ (5)
@@
Therefore $\mathbf{a}_{1:4}=(\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{u}_{1},\mathbf{e}_{3})$. Contrast this order with the more natural sort order that interleaves $\mathbf{u}_{1}$ as if it were an event at time $s_{1}-\delta=2$: $(\mathbf{e}_{1},\mathbf{u}_{1},\mathbf{e}_{2},\mathbf{e}_{3})$. Under the sort order, $\mathbf{u}_{1}$ appears before the first event $\mathbf{e}_{j}$ exceeding time $s_{1}-\delta$, i.e., the event $\mathbf{e}_{2}$ appearing at index $3$. This rule for placing $\mathbf{u}_{1}$ cannot be applied during autoregressive inference, as it requires us to place $\mathbf{u}_{1}$ at index $2$ with foreknowledge of the event $\mathbf{e}_{2}$ that has yet to be generated. Under the anticipatory order (Definition 3.1), $\mathbf{u}_{1}$ appears after the event $\mathbf{e}_{2}$ appearing at index $2$. This rule for placing $\mathbf{u}_{1}$ can be applied during the process of autoregressive inference: upon observing $\mathbf{e}_{2}$ at index $2$, we place $\mathbf{u}_{1}$ at index $3$. This distinction between the anticipatory order (which admits autoregressive inference) and the sort order (which does not) can be formalized using the concept of a stopping time, which we describe in Section 3.1.
@@ 3.2 Sparse Sequences
In this case, the anticipatory interleaving of $\mathbf{u}_{1}$ with $\mathbf{e}^{\prime}_{1:5}$ is $\mathbf{a}^{\prime}=(\mathbf{e}^{\prime}_{1},\mathbf{e}^{\prime}_{2},\mathbf{e}^{\prime}_{3},\mathbf{u}_{1},\mathbf{e}^{\prime}_{4},\mathbf{e}^{\prime}_{5})$. We anticipate the control $\mathbf{u}_{1}$ on time $s_{1}=4.5$ between the event $\mathbf{e}^{\prime}_{3}$ at time $3$ and the event $\mathbf{e}^{\prime}_{4}$ at time $4$. All the anticipatory models in this paper are trained with $\Delta^{*}=1$ second.
@@ 3.3 Training Anticipatory Models
We train anticipatory autoregressive models using standard maximum likelihood estimation of the sequences $\mathbf{a}_{1:N+K}$ (Definition 3.1). We tokenize these sequences according to the encoding described in Definition 2.2: $\mathbf{x}_{1:3(N+K)}\equiv\texttt{tokenize}(\mathbf{a}_{1:N+K})$. We follow a standard sequence packing procedure (see, e.g, Appendix B of Brown et al. (2020)) to construct training examples of fixed length $M$ (the model context) from variable-length sequences $\mathbf{x}_{1:3(N+K)}$, using a special event SEP as the sequence separator. We prepend each training example with a single global control code $\mathbf{z}\in\{0,1\}$ that indicates whether the example contains local controls $\mathbf{u}_{k}$; setting $\mathbf{z}=0$ (no controls) facilitates comparisons between anticipatory and autoregressive models. If the training example spans multiple sequences, then $\mathbf{z}$ describes the sequence preceding the first SEP event. We randomly shuffle and mini-batch training examples for stochastic gradient training.
@@
Dividing up a dataset into training examples of length $M$ introduces boundary effects. For general autoregressive models, this procedure results in an “early token curse,” whereby predictions early in a training example must be made with limited context (Press et al., 2021). For anticipatory autoregressive models, more subtle boundary effects arise. For an anticipation interval $\delta$, controls on first $\delta$ seconds of the training example do not appear in the context: these controls appear at the end of the previous training example. If $\delta$ is large relative to $M$ then the boundary effects become severe: many event predictions will be made without the relevant contextual controls, and vice-versa. This tempers the value of making predictions for all $M-1$ sequence indices, imposing a drag on training efficiency for large values $\delta$ relative to $M$. Like the early token curse, this effect is mitigated with larger contexts $M$ relative to $\delta$; in practice, the maximum practical anticipation interval $\delta$ is thus coupled with the context size of the model that we plan to train.
@@
$\mathbf{a}_{0}\leftarrow\texttt{SEP}$
  // A special sequence separator event
$i\leftarrow 1$
  // Index $i$ tracks position in the generated sequence
$k\leftarrow 0$
  // Index $k$ tracks position in the control sequence
@@
       $t\leftarrow\texttt{time}(\mathbf{a}_{i-1})$
        // Get the time $t$ of the previous event $\mathbf{a}_{i-1}$
       while $\emph{{time}}(\mathbf{u}_{k})\leq t+\delta$
        // While there are controls before time $t+\delta$
@@
             $\mathbf{a}_{i}\leftarrow\mathbf{u}_{k}$
              // Anticipate control $\mathbf{u}_{k}$ at index $i$
             $i\leftarrow i+1$
              // Advance to index $i+1$
             $k\leftarrow k+1$
              // Consume control $\mathbf{u}_{k}$
@@
      Sample $\mathbf{a}_{i}\sim p(\cdot|\mathbf{z},\mathbf{a}_{i-M-\text{Length}(\mathbf{z}):i-1})$
        // Sample an event from the model
       $i\leftarrow i+1$
        // Advance to index $i+1$